# Mamba-FSCIL: Dynamic Adaptation with Selective State Space Model for Few-Shot Class-Incremental Learning

### v1

@@ 4.2 Dual Selective SSM Projector
As illustrated in Fig. 2, a backbone network is trained in the base session to produce the intermediate feature ${\mathbf{F}}$, and will be frozen in the training of incremental sessions. To fully exploit the potential for dynamic adaptation, we develop a dual selective SSM projector consisting of three branches, i.e., an identity branch, and two selective SSM branches. The identity branch processes ${\mathbf{F}}$ through ${\bm{\mu}}^{\text{iden}}=p(\text{AvgPool}({\mathbf{F}}))$, where AvgPool is the average pooling function, $p$ is linear projection operation. In the selective SSM branch for the base session ($g^{\text{base}}$), ${\mathbf{F}}$ is first reshaped from a 2D feature map into a 1D vector to be processed by an MLP layer and is added with the position embedding as:
@@
where ${\mathbf{E}}^{\text{base}}\in\mathbb{R}^{L\times D^{\prime}}$ denotes the position embedding, $\hat{{\mathbf{F}}}^{\text{base}}\in\mathbb{R}^{N\times L\times D^{\prime}}$ is the output feature, $D^{\prime}$ denotes the output dimension, and $L=H\times W$ is the spatial locations. Then $\hat{{\mathbf{F}}}^{\text{base}}$ goes through a selective SSM projector, which splits the features into two information streams via linear projections $f_{x}$ and $f_{z}$, creating ${\mathbf{X}}^{\text{base}}=f_{x}(\hat{{\mathbf{F}}}^{\text{base}})$ and ${\mathbf{Z}}^{\text{base}}=f_{z}(\hat{{\mathbf{F}}}^{\text{base}})$. ${\mathbf{X}}^{\text{base}}$ is intended for scanning, while ${\mathbf{Z}}^{\text{base}}$ is used for gating the output of the scanned results. Specifically, the ${\mathbf{X}}^{\text{base}}$ stream is then enhanced through a group convolution function followed by a $\rm{SiLU}$ activation function:
@@
The enhanced features $\hat{{\mathbf{X}}}^{\text{base}}$ are then processed by the SS2D module defined in Eq. 5, enhancing discrimination and stability crucial for maintaining high accuracy within base classes. The output of the selective SSM branch ${\bm{\mu}}^{\text{base}}$ is obtained by the element-wise multiplication of the SS2D module’s output and ${\mathbf{Z}}^{\text{base}}$ (after applying the $\rm{SiLU}$ activation) with average pooling:
${\bm{\mu}}^{\text{base}}=\text{AvgPool}\left(\text{SS2D}(\hat{{\mathbf{X}}}^{\text{base}})\odot\text{SiLU}({\mathbf{Z}}^{\text{base}})\right),$ (8)
where $\odot$ denotes element-wise multiplication. The SS2D system matrices for the input features $\hat{{\mathbf{X}}}^{\text{base}}$ are calculated by linear projections as:
@@
where ${\mathbf{B}}\in\mathbb{R}^{N\times K\times D_{B}\times L}$, ${\mathbf{C}}\in\mathbb{R}^{N\times K\times D_{C}\times L}$, $\bm{\Delta}\in\mathbb{R}^{N\times K\times D_{\Delta}\times L}$ are data-dependent system matrices, with $K$ representing the number of scan directions in Eq. 5. Here, $D_{*}$ represents the dimensions of each parameter. Then, the final representation of the base session is obtained by combining the outputs of the identity branch and the selective SSM branch as:
@@
Concretely, in each incremental session $1\leq t\leq T$, the backbone network $f$, the identity branch $p$, and the selective SSM branch for the base session $g^{\text{base}}$ are frozen, and only the newly introduced $g^{\text{inc}}$ is learnable to adapt the model to novel classes. Following the widely adopted practice in FSCIL [7, 5, 3, 22, 70], we employ a memory module ${\mathcal{M}}^{(t)}$, which stores the mean intermediate feature ${\mathbf{F}}_{c}$ of each old class $c$ to alleviate forgetting when encountering novel classes:
@@
The intermediate features of the new class data, together with the stored mean features of the old classes, go through $g^{\text{inc}}$ to obtain the output:
${\bm{\mu}}^{\text{inc}}=\text{AvgPool}\left(\text{SS2D}(\hat{{\mathbf{X}}}^{\text{inc}})\odot\text{SiLU}({\mathbf{Z}}^{\text{inc}})\right),$ (12)
where ${\mathbf{X}}^{\text{inc}}$ and ${\mathbf{Z}}^{\text{inc}}$ are processed with the same architecture as $g^{\text{base}}$. The resulting final representation for incremental sessions is aggregated as:
@@ 4.3 Class-sensitive Selective Scan
In this section, we develop a class-sensitive selective scan mechanism that guides the dynamic adaptation of the selective SSM branch via two loss terms, i.e., the suppression loss and the separation loss.
Suppression Loss.
The feature vector ${\mathbf{Z}}^{\text{inc}}$ in  Eq. 12 plays a pivotal role in the element-wise multiplication with outputs from the SS2D module. By adjusting the norm of ${\mathbf{Z}}^{\text{inc}}$, we intend to manipulate the magnitude of the resulting feature vector ${\bm{\mu}}^{\text{inc}}$, which serves as a class-sensitive gating mechanism that controls the impact of new features with respect to base classes and incremental classes differently in the incremental session training.
  • For base classes, we reduce the norm of ${\mathbf{Z}}^{\text{inc}}$ of base classes to decrease ${\bm{\mu}}^{\text{inc}}$’s magnitude, thus minimizing feature shifts caused by the new branch $g^{\text{inc}}$ to maintain the stability of the features of the base class.
  • For novel classes, conversely, we increase the norm of ${\mathbf{Z}}^{\text{inc}}$ to enhance ${\bm{\mu}}^{\text{inc}}$’s magnitude, facilitating the learning of feature shifts that help the model adapt to new classes.
@@
In this way, the interference with base-class features brought by the newly added branch $g_{\text{inc}}$ is diminished by managing ${\mathbf{Z}}^{\text{inc}}$ strategically. In the ideal case, it is expected that the output of $g_{\text{inc}}$, i.e., ${\bm{\mu}}^{\text{inc}}$, could be zero such that the inference of base-class data after incremental session’s training (Eq. (13)) will be equivalent to the one after the base session’s training (Eq. (10)) to induce stability.
Separation Loss.
We further introduce a separation term that incurs class sensitivity in the input-dependent parameters (${\mathbf{B}},{\mathbf{C}},\bm{\Delta}$) in SS2D. For both base and novel classes, we calculate their average input-dependent parameters:
@@
where ${\mathbf{P}}$ represents any of ${\mathbf{B}}$, ${\mathbf{C}}$, or $\bm{\Delta}$ and $N_{\text{base}}$ and $N_{\text{novel}}$ denote the number of samples in base and novel classes, respectively. We then compute the cosine similarity of these average parameters between $\bar{{\mathbf{P}}}_{\text{b}}$ and $\bar{{\mathbf{P}}}_{\text{n}}$, aiming to minimize this similarity to maximize the separability of these selective scan parameters between base and novel classes. This process can be formulated as:
@@
As a result, the selective scan can be performed in distinct patterns between base and novel classes to induce plasticity in the incremental session’s training.
@@ 4.4 Optimization
Base Session Training.
During base session training ($t=0$), the backbone network $f$, the identity branch $p$, and the selective SSM branch for the base session $g^{\text{base}}$ are jointly optimized to establish a robust feature extractor. The objective during this phase is to minimize the classification loss on the extensive base session dataset:
@@
Incremental Sessions Training.
In subsequent sessions ($1\leq t\leq T$), as new classes are introduced, the optimization of the model focuses on seamlessly integrating these classes while minimizing disruption of previously learned knowledge. The training objective during these sessions is to adapt the model with new-class data and maintain the existing ability of old classes with the stored intermediate features in ${\mathcal{M}}^{(t)}$. Thus, the classification loss can be formulated as:
@@ 5 Experiments
Our experiments evaluate Mamba-FSCIL on the three wildly used benchmarks including miniImageNet [48], CIFAR-100 [24], and CUB-200 [60], comparing it against static and dynamic methods. We conduct ablation studies to assess the impact of our dual selective SSM projector and class-sensitive selective scan mechanism. We also provide an analysis of feature embeddings and compare our dynamic adaptation ability with Dynamic Support Networks (DSN) [68] to demonstrate efficiency in model complexity and superior performances across sessions.
5.1 Datasets and Implementation Details
Datasets. We employ three benchmark datasets. miniImageNet is a variant of ImageNet with 100 classes, each having 500 training and 100 testing images of $84\times 84$ pixels. CIFAR-100 has the same number of classes and images, and the image size is $32\times 32$. CUB-200 is a fine-grained classification dataset consisting of 11,788 images in 200 classes, with an image resolution of $224\times 224$ pixels. We follow the standard experimental settings in FSCIL [57, 76, 70]. For both miniImageNet and CIFAR-100, the base session contains 60 classes, and a 5-way 5-shot problem (5 classes and 5 images per class) is adopted for each of the 8 incremental sessions. For CUB-200, 100 classes are used in the base session, and there are 10 incremental sessions, each of which is a 10-way 5-shot.
Training Details. We employ the standard data preprocessing and augmentation techniques, including random resizing, flipping, and color jittering [57, 76, 42, 70]. We conduct training with a batch size of 512 for base sessions and 64 for incremental sessions, which includes new session data along with intermediate features stored in memory. For miniImageNet, we utilize ResNet12 [20] as the backbone and train for 500 epochs in the base session, followed by 100-170 iterations in each incremental session. The initial learning rate is 0.25 for the base session and 0.01 for the incremental sessions. For CIFAR-100, we utilize ResNet12 [20] as the backbone and train the base session for 200 epochs, followed by 200 iterations per incremental session. The initial learning rate is 0.25 for both base and incremental sessions. For CUB-200, we utilize ResNet18 [20] as backbone and train for 80 epochs in the base session, followed by 200-290 iterations in each incremental session. The initial learning rates are 0.02 for the backbone and 0.2 for the projector during the base session, and 0.05 for each incremental session. We apply a cosine annealing strategy for the learning rate and use SGD with a momentum of 0.009 and a weight decay of 0.0005 as the optimizer. The hyper-parameters $\alpha$ and $\beta$ are set between $50$ and $200$, $0.05$ and $0.5$, respectively.
Table 1: Performance of FSCIL in each session on miniImageNet and comparison with other studies. “Average Acc.” is the average accuracy of all sessions. “Final Improv.” calculates the improvement of our method in the last session over previous methods.
Methods Accuracy in each session (%) $\uparrow$ Average Final
@@ 5.1 Datasets and Implementation Details
IDLVQ [5] 64.77 59.87 55.93 52.62 49.88 47.55 44.83 43.14 41.84 51.16 +17.52
Self-promoted [82] 61.45 63.80 59.53 55.53 52.50 49.60 46.69 43.79 41.92 52.76 +17.44
SFbFSCIL [8] 61.40 59.80 54.20 51.69 49.45 48.00 45.20 43.80 42.10 50.63 +17.26
Data-free Replay [29] 71.84 67.12 63.21 59.77 57.01 53.95 51.55 49.52 48.21 58.02 +11.15
LIMIT [81] 72.32 68.47 64.30 60.78 57.95 55.07 52.70 50.72 49.19 59.06 +10.17
MetaFSCIL [9] 72.04 67.94 63.77 60.29 57.58 55.16 52.90 50.79 49.19 58.85 +10.17
C-FSCIL [22] 76.40 71.14 66.46 63.29 60.42 57.46 54.78 53.11 51.41 61.61 +7.95
Regularizer [3] 80.37 74.68 69.39 65.51 62.38 59.03 56.36 53.95 51.73 63.71 +7.63
CABD [78] 74.65 70.43 66.29 62.77 60.75 57.24 54.79 53.65 52.22 61.42 +7.14
ALICE [42] 80.60 70.60 67.40 64.50 62.50 60.00 57.80 56.80 55.70 63.99 +3.66
SAVC [52] 81.12 76.14 72.43 68.92 66.48 62.95 59.92 58.39 57.11 67.05 +2.25
NC-FSCIL [70] 84.02 76.80 72.00 67.83 66.35 64.04 61.46 59.54 58.31 67.82 +1.05
@@
TOPIC [57] 61.31 50.09 45.17 41.16 37.48 35.52 32.19 29.46 24.42 39.64 +34.94
DSN [68] 68.95 63.46 59.78 55.64 52.85 51.23 48.9 46.78 45.89 54.83 +13.47
CEC [76] 72.00 66.83 62.97 59.43 56.70 53.73 51.19 49.24 47.63 57.75 +11.73
@@
Table 2: Performance of FSCIL in each session on CIFAR-100 and comparison with other studies. “Average Acc.” is the average accuracy of all sessions. “Final Improv.” calculates the improvement of our method in the last session over previous methods.
Methods Accuracy in each session (%) $\uparrow$ Average Final

### v2

@@ 4.2 Dual Selective SSM Projector
As illustrated in Figure 2, a backbone network produces the intermediate feature ${\mathbf{F}}$. The dual selective SSM projector is composed of three branches: an identity branch, and two selective SSM branches. The identity branch processes ${\mathbf{F}}$ through ${\bm{\mu}}^{\text{iden}}=p(\text{AvgPool}({\mathbf{F}}))$, where AvgPool is the average pooling function, $p$ is linear projection operation. In the selective SSM branch for the base session ($g^{\text{base}}$), as shown in Fig. 2 (b), ${\mathbf{F}}$ is first reshaped from a 2D feature map into a sequence of flattened feature vectors to be processed by an MLP layer and is added with the position embedding as:
@@
where ${\mathbf{E}}^{\text{base}}\in\mathbb{R}^{L\times D^{\prime}}$ denotes the position embedding, $\hat{{\mathbf{F}}}^{\text{base}}\in\mathbb{R}^{N\times L\times D^{\prime}}$ is the output feature, $D^{\prime}$ denotes the output dimension, and $L=H\times W$ denotes the total number of spatial locations. Then $\hat{{\mathbf{F}}}^{\text{base}}$ goes through a selective SSM module, which splits the features into two information streams via linear projections $f_{x}$ and $f_{z}$, creating ${\mathbf{X}}^{\text{base}}=f_{x}(\hat{{\mathbf{F}}}^{\text{base}})$ and ${\mathbf{Z}}^{\text{base}}=f_{z}(\hat{{\mathbf{F}}}^{\text{base}})$. ${\mathbf{X}}^{\text{base}}$ is intended for scanning, while ${\mathbf{Z}}^{\text{base}}$ is used for gating the output of the scanned results. Specifically, the ${\mathbf{X}}^{\text{base}}$ stream is then enhanced through a group convolution function followed by a $\rm{SiLU}$ activation function:
@@
The enhanced features $\hat{{\mathbf{X}}}^{\text{base}}$ are then processed by the SS2D module defined in Eq. (5), enhancing discrimination and stability crucial for maintaining high accuracy within base classes. The output of the selective SSM branch ${\bm{\mu}}^{\text{base}}$ is obtained by elementwise multiplication of the output of the SS2D module and ${\mathbf{Z}}^{\text{base}}$ (after applying the $\rm{SiLU}$ activation) with average pooling:
${\bm{\mu}}^{\text{base}}=\text{AvgPool}\left(\text{SS2D}(\hat{{\mathbf{X}}}^{\text{base}})\otimes\text{SiLU}({\mathbf{Z}}^{\text{base}})\right),$ (8)
where $\otimes$ denotes elementwise multiplication. The SS2D system matrices for the input features $\hat{{\mathbf{X}}}^{\text{base}}$ are calculated by linear projections as:
@@
where ${\mathbf{B}}\in\mathbb{R}^{N\times K\times D_{B}\times L}$, ${\mathbf{C}}\in\mathbb{R}^{N\times K\times D_{C}\times L}$, $\bm{\Delta}\in\mathbb{R}^{N\times K\times D_{\Delta}\times L}$ are data-dependent system matrices, with $K$ representing the number of scan directions in Eq. (5). Here, $D_{*}$ represents the dimensions of each parameter. Then, the final representation of the base session is obtained by combining the outputs of the identity branch and the selective SSM branch as:
@@
Concretely, in each incremental session $1\leq t\leq T$, the backbone network $f$, the identity branch $p$, and the selective SSM branch for the base session $g^{\text{base}}$ are frozen, and only the newly introduced $g^{\text{inc}}$ is learnable to adapt the model to novel classes. Following the widely adopted practice in FSCIL [10, 11, 12, 16, 19, 82], we employ a memory module ${\mathcal{M}}^{(t)}$, which stores the mean intermediate feature ${\mathbf{F}}_{c}$ of each old class $c$ to alleviate forgetting when encountering novel classes:
@@
The intermediate features of the new class data, along with the stored mean features of the old classes, are processed through $p$ and $g^{\text{base}}$ to produce ${\bm{\mu}}^{\text{iden}}$ and ${\bm{\mu}}^{\text{base}}$ respectively, using the same processes as in the base session. Additionally, these features are processed through $g^{\text{inc}}$ to generate the output:
${\bm{\mu}}^{\text{inc}}=\text{AvgPool}\left(\text{SS2D}(\hat{{\mathbf{X}}}^{\text{inc}})\otimes\text{SiLU}({\mathbf{Z}}^{\text{inc}})\right),$ (12)
where $\hat{{\mathbf{X}}}^{\text{inc}}$ and ${\mathbf{Z}}^{\text{inc}}$ are processed using the same architecture as the base branch $g^{\text{base}}$, but with different parameters. The resulting final representation for incremental sessions is aggregated as:
@@ 4.3 Class-sensitive Selective Scan
In this section, we develop a class-sensitive selective scan mechanism, as illustrated in Fig. 2 (c) that guides the dynamic adaptation of the incremental selective SSM branch $g^{\text{inc}}$ by two loss terms: suppression loss (${\mathcal{L}}_{\text{supp}}$) and separation loss (${\mathcal{L}}_{\text{sep}}$).
Suppression Loss (${\mathcal{L}}_{\text{supp}}$). This loss focuses on adjusting the norm of the feature vector $\mathbf{Z}^{\text{inc}}$, depicted in the top part of Fig. 2 (c). The diagram shows the distribution of features for the base (blue dots) and novel (orange dots) classes, with dashed lines representing the intention of modifying feature norms. In Eq. (12), the feature vector ${\mathbf{Z}}^{\text{inc}}$ plays a pivotal role in the elementwise multiplication with outputs from the SS2D module. By adjusting the norm of ${\mathbf{Z}}^{\text{inc}}$, we intend to manipulate the magnitude of the resulting feature vector ${\bm{\mu}}^{\text{inc}}$, which serves as a class-sensitive gating mechanism that controls the impact of new features with respect to base classes and incremental classes differently during incremental session training:
  • For base classes, the norm of ${\mathbf{Z}}^{\text{inc}}$ of base classes is suppressed (red dashed lines) to decrease ${\bm{\mu}}^{\text{inc}}$’s magnitude, thus minimizing feature shifts caused by the new branch $g^{\text{inc}}$ to maintain the stability of the features of the base classes.
  • For novel classes, conversely, the norm of ${\mathbf{Z}}^{\text{inc}}$ is increased (green dashed lines) to enhance ${\bm{\mu}}^{\text{inc}}$’s magnitude, facilitating the learning of feature shifts that help the model adapt to new classes.
@@
By differentially managing the feature magnitudes for base and novel classes through suppression loss, the interference with base-class features brought by the newly added branch $g_{\text{inc}}$ is diminished, ensuring that new class integration does not destabilize the model’s existing knowledge. In the ideal case, it is expected that the output of $g^{\text{inc}}$, i.e., ${\bm{\mu}}^{\text{inc}}$, could be zero such that the inference of base-class data after incremental session’s training (Eq. (13)) will be equivalent to the one after the base session’s training (Eq. (10)) to induce stability.
Separation Loss (${\mathcal{L}}_{\text{sep}}$). The separation term is designed to enhance class sensitivity in the input-dependent parameters (${\mathbf{B}},{\mathbf{C}},\bm{\Delta}$) within the SS2D module, ensuring that base and novel classes are effectively distinguished. As illustrated in the lower part of Fig. 2 (c), the separation mechanism ”pushes” the averaged input-dependent parameters associated with novel classes (${\mathbf{B}}_{n},{\mathbf{C}}_{n},\bm{\Delta}_{n}$) away from those of the base classes (${\mathbf{B}}_{b},{\mathbf{C}}_{b},\bm{\Delta}_{b}$), thereby maximizing the distinctiveness between these classes.
Specifically, for both base and novel classes, we calculate their average input-dependent parameters:
@@
where $\bar{{\mathbf{P}}}_{\text{b}}$ and $\bar{{\mathbf{P}}}_{\text{n}}$ represent the aggregated parameters for the base and novel classes, respectively. The symbol ${\mathbf{P}}$ represents one of ${\mathbf{B}}$, ${\mathbf{C}}$, or $\bm{\Delta}$. The terms $N_{\text{base}}$ and $N_{\text{novel}}$ denote the number of samples in base and novel classes in a mini-batth. To maximize separability, we then compute and minimize the cosine similarity of these averaged parameters between base and novel classes ($\bar{{\mathbf{P}}}_{\text{b}}$ and $\bar{{\mathbf{P}}}_{\text{n}}$), which can be formulated as:
@@
By minimizing ${\mathcal{L}}_{\text{sep}}$, the selective scan can be performed in distinct patterns between base and novel classes to induce plasticity during incremental learning.
@@ 4.4 Optimization
Base Session Training. During base session training ($t=0$), the backbone network $f$, the identity branch $p$, and the selective SSM branch for the base session $g^{\text{base}}$ are jointly optimized to establish a robust feature extractor. The objective during this phase is to minimize the classification loss on the extensive base session dataset:
@@
Incremental Sessions Training. In subsequent sessions ($1\leq t\leq T$), as new classes are introduced, the optimization of the model focuses on seamlessly integrating these classes while minimizing disruption of previously learned knowledge. The training objective during these sessions is to adapt the model with new-class data and maintain the existing ability of old classes with the stored intermediate features in ${\mathcal{M}}^{(t)}$. Thus, the classification loss can be formulated as:
@@ 5 Experiments
Our experiments evaluate Mamba-FSCIL on the three widely used benchmarks, including miniImageNet [112], CIFAR-100 [113], and CUB-200 [114], comparing it with static and dynamic methods. We conducted ablation studies to analyze the contributions of key components, including the dual selective SSM projector and the class-sensitive selective scan mechanism, as well as the effects of suppression and separation losses on feature alignment and class separability. We also examined the impact of freezing versus training the identity and selective SSM branches, the sensitivity to hyper-parameters, and the influence of varying training samples (shots) and scan paths ($K$). Additionally, we compared different projection architectures and evaluated Mamba-FSCIL against Dynamic Support Networks (DSN) [23] to demonstrate its efficiency in managing model complexity while achieving superior performance. Finally, we visualized activation maps to further validate the effectiveness of our approach and provided an analysis of feature embeddings.
TABLE I: FSCIL performance across sessions on miniImageNet compared with other methods. ”Average Acc.” is the average accuracy across all sessions. ”Final Improv.” denotes the accuracy improvement of our method in the last session compared to previous methods.
Methods Venue Accuracy in each session (%) $\uparrow$ Average Final
@@
IDLVQ [11] ICLR 2021 64.77 59.87 55.93 52.62 49.88 47.55 44.83 43.14 41.84 51.16 +17.52
Self-promoted [86] CVPR 2021 61.45 63.80 59.53 55.53 52.50 49.60 46.69 43.79 41.92 52.76 +17.44
SFbFSCIL [80] ICCV 2021 61.40 59.80 54.20 51.69 49.45 48.00 45.20 43.80 42.10 50.63 +17.26
Data-free Replay [6] ECCV 2022 71.84 67.12 63.21 59.77 57.01 53.95 51.55 49.52 48.21 58.02 +11.15
LIMIT [18] TPAMI 2022 72.32 68.47 64.30 60.78 57.95 55.07 52.70 50.72 49.19 59.06 +10.17
MetaFSCIL [17] CVPR 2022 72.04 67.94 63.77 60.29 57.58 55.16 52.90 50.79 49.19 58.85 +10.17
FACT [3] CVPR 2022 72.56 69.63 66.38 62.77 60.6 57.33 54.34 52.16 50.49 60.70 +8.87
C-FSCIL [16] CVPR 2022 76.40 71.14 66.46 63.29 60.42 57.46 54.78 53.11 51.41 61.61 +7.95
Regularizer [12] ICLR 2022 80.37 74.68 69.39 65.51 62.38 59.03 56.36 53.95 51.73 63.71 +7.63
TEEN [87] NeurIPS 2023 73.53 70.55 66.37 63.23 60.53 57.95 55.24 53.44 52.08 61.44 +7.28
CABD [88] CVPR 2023 74.65 70.43 66.29 62.77 60.75 57.24 54.79 53.65 52.22 61.42 +7.14
ALICE [8] ECCV 2022 80.60 70.60 67.40 64.50 62.50 60.00 57.80 56.80 55.70 63.99 +3.66
NC-FSCIL [19] ICLR 2023 84.02 76.80 72.00 67.83 66.35 64.04 61.46 59.54 58.31 67.82 +1.05
@@
TOPIC [1] CVPR 2020 61.31 50.09 45.17 41.16 37.48 35.52 32.19 29.46 24.42 39.64 +34.94
LEC-Net [21] arXiv 2022 61.31 35.37 36.66 38.59 33.90 35.89 36.12 32.97 30.55 37.92 +29.69
DSN [23] TPAMI 2022 68.95 63.46 59.78 55.64 52.85 51.23 48.9 46.78 45.89 54.83 +13.47
CEC [2] CVPR 2021 72.00 66.83 62.97 59.43 56.70 53.73 51.19 49.24 47.63 57.75 +11.73
FeSSSS [89] CVPR 2022 81.50 77.04 72.92 69.56 67.27 64.34 62.07 60.55 58.87 68.23 +0.49
@@
5.1 Implementation Details
Datasets. We employed three benchmark datasets. miniImageNet is a variant of ImageNet with 100 classes, each having 500 training and 100 testing images of $84\times 84$ pixels. CIFAR-100 has the same number of classes and images, and the image size is $32\times 32$. CUB-200 is a fine-grained classification dataset consisting of 11,788 images in 200 classes, with an image resolution of $224\times 224$. We followed the standard experimental settings in FSCIL [1, 2, 19]. For miniImageNet and CIFAR-100, the base session includes 60 classes, followed by 8 incremental sessions, each with a 5-way 5-shot setting (5 classes and 5 images per class). For CUB-200, the base session comprises 100 classes, with 10 incremental sessions, each following a 10-way 5-shot setting.
Training Details. We employed the standard data preprocessing and augmentation techniques, including random resizing, flipping, and color jittering [1, 2, 8, 19]. We conducted training with a batch size of 512 for base sessions and 64 for incremental sessions, which includes new session data along with intermediate features stored in memory.
For miniImageNet, we utilized ResNet12 [115] as the backbone and train for 500 epochs in the base session, followed by 100-170 iterations in each incremental session. The initial learning rate is 0.25 for the base session and 0.01 for the incremental sessions.
For CIFAR-100, we utilized ResNet12 [115] as the backbone and trained the base session for 200 epochs, followed by 200 iterations per incremental session. The initial learning rate is 0.25 for both the base and incremental sessions.
For CUB-200, we employed two different architectures. In Table III, ResNet18 [115] serves as the backbone, where we trained the model for 80 epochs in the base session and then proceeded with 200-290 iterations for each incremental session. In Table IV, we utilized the Swin Transformer Tiny [30] with an input resolution of $384\times 384$, we trained the model for 80 epochs during the base session followed by 600 iterations per incremental session. Initial learning rates are set at 0.02 for the backbone and 0.2 for the projector during the base session, and are reduced to 0.05 for the incremental sessions.
We apply a cosine annealing strategy for the learning rate and use SGD with a momentum of 0.009 and a weight decay of 0.0005 as the optimizer. The hyper-parameters $\alpha$ and $\beta$ are set between $50$ and $200$, $0.05$ and $0.5$, respectively. The dimensions of the input-dependent parameters in Eq. (9) are set as $D_{B}=D_{C}=D_{\Delta}=256$. The dimensions $D^{\prime}$ in Eq. (6) are set to $512$, $512$, $1024$ for CIFAR-100, CUB-200, and miniImageNet datasets, respectively. All experiments are conducted using 8 NVIDIA A100 GPUs.
Zero Initialization. In our incremental learning process, we fixed the identity branch and selective SSM branch from the base session $g^{\text{base}}$ and introduced a new selective SSM branch $g^{\text{inc}}$ for the incremental sessions. To prevent disruption of pre-existing feature representations, we applied zero initialization to $g^{\text{inc}}$ by setting the projection parameters of $f_{z}$ for the gating features ${\mathbf{Z}}^{\text{inc}}$ to zero, ensuring that its initial outputs (${\bm{\mu}}^{\text{inc}}$) are also zero. This approach ensures that $g^{\text{inc}}$ begins without affecting the model’s learned features from the base session. As training progresses, $g^{\text{inc}}$ is gradually optimized, allowing the model to adapt to new classes by learning feature shifts while preserving the knowledge acquired during the base session.
5.2 Comparison with the State-of-the-art Methods
TABLE II: FSCIL performance across sessions on CIFAR-100 compared with other methods. ”Average Acc.” is the average accuracy across all sessions. ”Final Improv.” denotes the accuracy improvement of our method in the last session compared to previous methods.
Methods Venue Accuracy in each session (%) $\uparrow$ Average Final